OpenStudy (anonymous):

verify each identity: cos(alpha+beta)cos(alpha-beta)=cos^2beta-sin^2alpha

OpenStudy (noorik.):

the laws: cos(alpha+beta) = cos(alpha)*cos(beta) - sin(alpha)*sin(beta) cos(alpha-beta) = cos(alpha)*cos(beta) + sin(alpha)*sin(beta) they would help you

OpenStudy (anonymous):

i used those but i got stuck :(

OpenStudy (noorik.):

why? cos(alpha+beta)cos(alpha-beta) = [cos(alpha)*cos(beta) - sin(alpha)*sin(beta)] * [cos(alpha)*cos(beta) + sin(alpha)*sin(beta)] = [cos(alpha)*cos(beta)]^2 - [sin(alpha)*sin(beta)]^2 = [cos^2(alpha)*cos^2(beta)] - [sin^2(alpha)*sin^2(beta)] since cos^2(alpha) = 1- sin^2(alpha) and sin^2(beta) = 1-cos^2(beta) , then = [cos^2(beta)-sin^2(alpha)*cos^2(beta)] - [sin^2(alpha)-sin^2(alpha)*cos^2(beta)] then sin^2(alpha)*cos^2(beta) - sin^2(alpha)*cos^2(beta) = 0 so = cos^2(beta) - sin^2(alpha)

OpenStudy (anonymous):

ohhh ok :) thanks! i made a mistake in one of my steps

OpenStudy (noorik.):

u r welcome